OpenStudy (anonymous):

Solve 4x^2=4 What are the solutions? what are the intercepts of f(x)=4x^2-8?

OpenStudy (anonymous):

First one: 4x^2=4 now divide both sides by 4 and we get: x^2=1 now take the square root of both sides and we get: x=plus or minus 1

OpenStudy (anonymous):

Second: is the same as the first, x intercepts can be solution Do, the same process as i did for the first one, you should get : plus or minus \[\sqrt{2}\]

OpenStudy (anonymous):

so (1+sqrt2),(1-sqrt2)? is this correct?

OpenStudy (anonymous):

for which one?

OpenStudy (anonymous):

the second one is only plus or minus sqrt 2

OpenStudy (anonymous):

The top? I'm confused lol

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

4x^2=4 is only plus or minus 1

OpenStudy (anonymous):

f(x)=4x^2-8 is only plus or minus sqrt 2

OpenStudy (anonymous):

so -1,2?

OpenStudy (anonymous):

for the first part?

OpenStudy (anonymous):

where die you get a two from?

OpenStudy (anonymous):

the intercepts for the first part are: +1 and -1

OpenStudy (anonymous):

lol told u I was confused I see now

OpenStudy (anonymous):

the intercepts for the second one are : +sqrt2 and -sqrt2

OpenStudy (anonymous):

got it!